OpenStudy (anonymous):

Determine the zeros of f(x) = x4 – x3 + 7x2 – 9x – 18.

OpenStudy (anonymous):

I don't understand how to do this one @phi

OpenStudy (phi):

± factors of 18: ± 1,2,3,6,9,18 so try and see if you can find a root

OpenStudy (anonymous):

2

OpenStudy (phi):

now do synthetic division

OpenStudy (anonymous):

x^3+x^2+9x+9=0

OpenStudy (phi):

yes, now try to find a root for this using ±1,3,9

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

sy

OpenStudy (phi):

ok, now syn div on the cubic

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x=+/-3

OpenStudy (phi):

almost you should have gotten x^2 +9 = 0 did you ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

*x=+/-3i sorry

OpenStudy (phi):

then x^2= -9 now as you know, there is no real number that when you square it is *negative*

OpenStudy (phi):

yes, looks good